length of a dog run is 9 feet more than twice the width and the perimeter is 34 feet, what is the length and width
Answer by tommyt3rd(5050)

1) l=2w+9
2) 34=2(l+w)

17=l+w
17=2w+9+w
17-9=3w
w=8/3

so
l=
2*8/3+9=
16/3+27/3=
43/3